Bearclaws72:

Which expression is equivalent to one over fivem − 20? one over five(m − 4) one over five(m − 100) 5(m − 4) 5(m − 100)

Vocaloid:

same thing as before (1/5)(m-20) = (1/5)m - (1/5)20 = ?

Bearclaws72:

-3.8

Vocaloid:

one last time, you can't combine like terms

Vocaloid:

(1/5)20 = ?

Bearclaws72:

4

Vocaloid:

good so your answer is (1/5)m - 4

Bearclaws72:

k

Bearclaws72:

ill tell you what i get

Vocaloid:

wait

Bearclaws72:

yeah

Vocaloid:

is the original problem 1/5(m-20) or (1/5)m - 20

Bearclaws72:

yes

Vocaloid:

no, which one is it?

Vocaloid:

is the (m-20) inside parentheses or not

Bearclaws72:

no

Bearclaws72:

there not

Vocaloid:

then I think it's the second answer

Bearclaws72:

ok

Bearclaws72:

yes i got a 100 thx so much